OpenStudy (anonymous):

:cos alpha cos(60degree-alpha)cos(60degree+alpha)=1/4cos 3 alpha

OpenStudy (dumbcow):

are they being multiplied together?

OpenStudy (anonymous):

yes.....

OpenStudy (dumbcow):

angle sum and difference formulas for cos \[\cos(a+b) = \cos a*\cos b - \sin a*\sin b\] \[\cos(a-b) = \cos a*\cos b +\sin a*\sin b\] cos(60) = 1/2

OpenStudy (anonymous):

and aftr dat??

OpenStudy (dumbcow):

\[\cos(a+60) = (1/2)\cos a - (\sqrt{3}/2)\sin a\] \[\cos(a-60) = (1/2)\cos a + (\sqrt{3}/2)\sin a\] \[\cos(a+60)*\cos(a-60) = (1/4)\cos^{2} a - (3/4)\sin^{2} a\]

OpenStudy (dumbcow):

multiply by cos(a) \[\rightarrow (1/4)\cos^{3} a - (3/4)\cos a*\sin^{2} a\] sin^2 = 1 - cos^2 \[\rightarrow (1/4)\cos^{3} a - (3/4)\cos a*(1-\cos^{2} a) = (1/4)\cos^{3} a +(3/4)\cos^{3} a -(3/4)\cos a\] \[= \cos^{3} a - (3/4)\cos a\]

OpenStudy (dumbcow):

ok yeah this equals (1/4)cos(3a) \[\cos 3a = \cos 2a*\cos a - \sin 2a*\sin a\] \[= \cos a(2\cos^{2} a -1) - 2\cos a*\sin^{2} a\] \[= (2\cos^{3} a-\cos a) - 2\cos a(1-\cos^{2} a)\] \[= 4\cos^{3} a - 3\cos a\]